OpenStudy (asfcahgcax):

Factorise and simplify the following without initially expanding the brackets. 1.(x+2)^2-(x+3)^2 2.(y-7)^2-(y+4)^2 Answers 1.-(2x+5) 2.-11(2y-3) How can I get the answer?????

OpenStudy (skullpatrol):

Study these examples http://openstudy.com/study#/updates/57a3530fe4b0abc5c3d1dc2c

OpenStudy (skullpatrol):

Welcome to open study :D

OpenStudy (welshfella):

a^2 - b^2 = (a - b)(a + b) now replace a by x+2 and b by x+3 what do you get?

OpenStudy (welshfella):

replace the a and b in the right side

OpenStudy (asfcahgcax):

I keep getting (5)(2x+5)

OpenStudy (welshfella):

(x + 2 - (x + 3)] (x+ 2+x + 3) your mistake is not putting the x + 3 in parenthesis in the first part - You must subtract the WHOLE of (x + 3) so you put it in parentheses. Then you distribute the - over the parentheses and the + 3 becomes -3. = (x + 2 - x - 3)(2x + 5) = -1 (2x + 5) = - (2x + 5)

OpenStudy (welshfella):

get it?

OpenStudy (asfcahgcax):

oh I get it now Thanks a lot!!!

OpenStudy (welshfella):

yw